Nov 19, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Update BCAL clusterizer to improve agreement between simulated and generated MC (Mark)
  2. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt).
  3. Change F250 simulation code to match firmware and add flexibility for new algorithms (Mike)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Preparations
    1. FCAL Update
    2. BCAL Update
  4. Reconstruction and Simulation
  5. Calibration
    1. Calibration Challenge
  6. Any other business


Attendees: Elton, Mark, Adesh, David (JLab); Sean (NW); Zisis, Ahmed, Noemi, Tegan, Andrei (UofR); Will, Mike (CMU); Matt, Jon (IU)

  1. Announcements
  2. Action Items
  3. Run Preparations
    1. FCAL Update (Adesh)
      • Matt and Adesh replaced 52 bases last week that showed HV=0 and reported no signals in pulser data
      • Monitoring with EPICs since then shows about 6 pjs with HV=0 and checked no response with pulser data
      • Log Entry 3360930 (Jon) Four bases returned to IU found to have corrupt firmware. Reprogrammed and seem ok. Adesh was confused since these bases had a history of HV=0. They were also reprogrammed at JLab before return. Jon: After reprogrammed at IU they can be set and readback 1500 V correctly. They will be monitored for a week and checked before return to JLab.
    2. Gain calibration
      • Checked pedestals in runs with mode 7 and mode 8. All within +/- 1 count, which will be entered into ccdb
      • Log Entry 3360955 Pi0 calibration
      • Different calibration constants for mode 7 and mode 8. They differ by about 20%.
      • Mode 7: sigma~8.6 MeV, mean=134.8 MeV
      • Mode 8: sigma~11 MeV, mean =135.1 MeV.
      • Discrepancy in calibrations between mode 7 and mode 8 may be hinted by the different int/peak value mode7=6.5 (Run 2931) and mode 8=7.8 (Run 3079). This difference might be due to different Nsa+Nsb values for firmware and emulation. Adesh will check into this.
      • Next step is to check in the calibration constants into the ccdb for the two data modes
      • The HV on the tubes will be adjusted next week to decrease the variation in pmt gains.
      • Paul's new GUI for scalers will be used to check for hot channels.
    3. BCAL Update
  4. Reconstruction and Simulation
    1. Updates from Regina (Zisis)
      • Noemi continues to investigate the form of time-walk function comparing various options, which require adding columns to the ccdb entry. She will present comparison at the next meeting.
      • Tegan and Andrei are revisiting the sampling fraction data generated almost 2 years ago by Andrei and Irina. They are considering the best way to incorporate into hdgeant
    2. FCAL Hadron reconstruction Log Entry 3360676 (Elton)
      • Investigating the response of FCAL to hadrons. The energy deposition in the data exhibit a prominent minimum ionizing peak, a shoulder at higher energies thought to be due to Cherenkov light generated in the light guide, and a broad spectrum beginning at zero due to interacting pions. A rough efficiency for detecting a minimum energy in the FCAL per incident pion is shown
    3. FCAL MC nonlinear corrections Log Entry 3361053 (Elton)
      • The MC and data reconstruction use very different set of non-linear corrections. For the data, the constants effectively eliminate the correction. For the MC, the corrections vary from 0.5 to 2 for energies below 6 GeV.
      • We need to rethink how to incorporate both a scale factor into the MC as well as any non-linear corrections. We need to treat the data and MC on an equal footing.
    4. Timing studies (Mike)
      • Updated timing calibration tables for BCAL have been completed, but still need to be added to ccdb.
  5. Calibration
    1. Calibration Challenge
      • Updates are trickling in. Will likely postpone tomorrow's meeting till early next week.
      • Would like to include Mike's work. Will has also promised to provide a plugin for the BCAL gain calibration
      • Sean: Asked about omega->pi0 g events for calibration. Jon has been working on the signal MC for this. Matt: The angular resolution of the missing photon is about 1 degree, which might limit how well this reaction can be used for efficiency studies.
      • Sean: Are gain uncertainties available? Adesh/Will: No. Matt: Uncertainties should probably be included at the shower level. Elton:The uncertainties are driven by the calorimeter resolutions. We should probably get a parameterization of the resolution (which may also be a function of position) that can be used for input for the software to propagate through to higher level quantities.
  6. Any other business